# On Quantum Random Walks in Biomolecular Networks

### v1

@@ II-A Biomolecular Network
Then, the degree matrix $D\in\mathbb{R}^{n\times n}$ is defined as $D_{ii}=\sum_{j=1}^{n}A_{ij}\text{ }\forall\text{ }i\neq j$, and $D_{ij}=0$. The Laplacian matrix $L\in\mathbb{R}^{n\times n}$ of graph $G$ is defined as $L=D-A$.
@@ II-E Continuous-time Quantum Random Walk
[…]
@@
When considering a specific disease in the disease gene prioritization task, we do not need the whole matrix exponential but rather its action on the initial state vector |$\ket{\psi(0)}$. Consequently, calculating the scores can be an efficient process [21].
@@ II-F Discrete-time Quantum Random Walk
|$\ket{\psi(t)}=\sum_{j=1}^{n}\sum_{k=1}^{d_{j}}\psi_{j,k}(t)\ket{j\rightarrow k}$ (7)
where $n$ is defined as the number of vertices, as above and |$\ket{j\rightarrow k}$ is the state residing on vertex $j$ and is about to hop on to vertex $k$, and $d_{j}$ is the degree of the vertex $j$. The transition probability in this case is defined as
@@ II-G Performance Metrics
$P@K=\frac{\text{Number of relevant genes in the first }K\text{ ranks}}{K}$

### v2

@@ II-A Biomolecular Network
Then, the degree matrix $D\in\mathbb{R}^{n\times n}$ is defined as $D_{ii}=\sum_{j=1}^{n}A_{ij}\text{ }\forall\text{ }i\neq j$, and $D_{ij}=0$. The Laplacian matrix $L\in\mathbb{R}^{n\times n}$ of graph $G$ is defined as L=DAL=D-A.
@@ II-E Continuous-time Quantum Random Walk
A classical continuous-time random walk (CTRW) on a graph $G$ is defined by a walker that moves randomly between the vertices $V$ of the graph with random waiting times at each node governed by a continuous probability distribution. For a graph $G$ with $n$ vertices ($|V|=n)$, the classical walker uses a continuous-time Markovian process with a vector of probabilities $p$ to make the jump to a neighboring vertex, guided by
@@
[…]
@@ II-F Discrete-time Quantum Random Walk
[…]
@@ II-G Performance Metrics
[…]